OpenStudy (anonymous):

Find the area of the region under y=e^x and above y=1 for 0 less than equal to x less than equal to 5.

OpenStudy (anonymous):

Calculus huh? k, you need the definite integral of the function on top minus the function on bottom from 0 to 5 so...\[\int\limits_{0}^{5}e^x-1\]

OpenStudy (anonymous):

Do you take the antiderivatives of those?

OpenStudy (amistre64):

make sure the function dont cross at some point; in this case they remain separate tho

OpenStudy (amistre64):

and yes, integration is the word you use most after learning what an antiderivative is

OpenStudy (anonymous):

so it would be e^5-5

OpenStudy (amistre64):

-e^0

OpenStudy (amistre64):

(e^5-5)-(e^0-0) = e^5-5-1 = e^5-6

OpenStudy (anonymous):

THANKS!!!